# SCILLA: SurfaCe Implicit Learning for Large urban Area, a volumetric hybrid solution

### v3

@@ 9 Additional implementation details
We use the poses provided by the datasets, except for Waymo, where we recompute the vehicle trajectory and sensor calibration with MOISST [13] due to inaccuracies in the provided data. The overall loss we use to optimize SCILLA is defined as follows:
@@ 10.1 Additional qualitative results
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
@@
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
@@
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
@@
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
@@
Figure 8: Qualitative experiments results on 8(a) KITTI-360s [20], 8(b) Pandaset [41], 8(c) nuScenes [3] and 8(d) Waymo Open Dataset [32]. RGB ground-truth images shown for reference of the scene layout. Complex scene geometries reconstructed in the mesh are highlighted in red squares. We compare our generated SDF mesh to COLMAP, OpenMVS and StreetSurf [11] meshes.
We report in Fig.8 additional qualitative results on nuScenes [3] and Waymo Open Dataset [32]. We find that SCILLA reconstructs higher-quality surfaces compared to StreetSurf and can recover many scene details (see highlighted red-boxes on the figures).
@@ 10.2 Ablation study
We ablate the effect of random sample attribution and our proposed probability-based samples attribution. The qualitative results at various training steps are presented in  Fig.9. The results demonstrate that SCILLA samples attribution strategy initially learns the coarse geometry of the scene during early training stages. While random sample attribution can approximate an accurate SDF representation by the end of the hybrid stage, it results in an incomplete mesh compared to the mesh generated using probability-based sample attribution.
@@ 11 Applications
11.1 Tanks And Temples dataset
In this section, we report SCILLA’s qualitative results on Tanks and Temples dataset [18]. The results reported in Fig.  10 demonstrate that SCILLA is not restricted to driving datasets and urban areas. SCILLA can represent large, detailed, indoor and outdoor sequences from Tanks And Temples. We report the results of SCILLA with and without the hybrid volumetric stages. Our observation suggests that the transition from the volumetric to surface representation enables to recover various scene details. In addition, SCILLA’s training time takes less than an hour, compared to other state-of-the-art indoor and outdoor landmarks reconstruction methods, such as Neuralengelo [19] which requires 24h of training.
TanksAndTemples [18]
RGB Refer to caption Refer to caption Refer to caption Refer to caption
w/o hybrid Refer to caption Refer to caption Refer to caption Refer to caption
SCILLA (ours) Refer to caption Refer to caption Refer to caption Refer to caption
(a) Auditorium (b) Temple (c) Truck (d) Meetingroom
Figure 10: Qualitative experiments on TanksAndTemples dataset [18]. We report the RGB ground-truth image. In a different point-of-view than the GT RGB, we compare the extracted mesh from the SDF from both SCILLA, and SCILLA without the hybrid stages.
@@ 11.2 Textured mesh
Due to the high-quality of SCILLA’s reconstructed surfaces, we can leverage modern Multi-View Stereo (MVS) tools like OpenMVS [4] to produce detailed and colorized representations of driving sequences. As shown in Fig. 11, we find that SCILLA’s textured mesh is more complete and accurate compared to StreetSurf’s textured mesh.
SCILLA (ours) Refer to caption Refer to caption
@@
(c) SCILLA (ours) (d) StreetSurf [11]
Figure 12: Failure case of SCILLA for the sequence 31 from KITTI-360 [20]. We report the ground-truth RGB image and the predicted sky mask. In a different point-of-view than the GT RGB, we compare our generated SDF mesh to StreetSurf’s mesh.
@@
(c) SCILLA (ours) (d) StreetSurf [11]
Figure 13: Failure case of SCILLA for the sequence 23 from Pandaset [41]. We report the ground-truth RGB image and the wide part of the sequence highlighted. In a different point-of-view than the GT RGB, we compare our generated SDF mesh to StreetSurf’s mesh.
@@
(c) SCILLA (ours) (d) StreetSurf [11]
Figure 14: Failure case of SCILLA for the sequence 664 from nuScenes [3]. We report the ground-truth RGB image and the monocular normal. In a different point-of-view than the GT RGB we compare our generated SDF mesh to StreetSurf’s mesh.
@@ 12 Limitations
Although SCILLA’s reconstructed surfaces are highly detailed and accurate, we find that our method can fail in three distinct scenarios:
  • Disentangling fine details from the sky: unlike StreetSurf [11], which models close range, far range, and sky separately, SCILLA separates only the sky from the other scene’s modeling. However, SCILLA may struggle to distinguish fine details from the sky, particularly in cases where objects are thin, as shown in Fig. 12.
  • Fine details in wide sequences: SCILLA sometimes fails in accurately reconstructing scene details in wide and open sequences, such as scene 23 from Pandaset reported in Fig. 13.
  • Inaccurate road reconstruction: for sequence 664 from nuScenes (see Fig. 14), SCILLA faces challenges in reconstructing the road due to inaccuracies in the monocular normal prediction from Omnidata [8]. In comparison, StreetSurf [11] demonstrates more accurate road reconstruction, attributed to its road-surface initialization.

### v4

@@ 9 Additional implementation details
We use the poses provided by the datasets, except for Waymo, where we recompute the vehicle trajectory and sensor calibration with MOISST [14] due to inaccuracies in the provided data. The overall loss we use to optimize ViiNeuS is defined as follows:
@@ 10 Additional results
10.1 Additional quantitative results
KITTI-360 [21] Pandaset [43] Waymo [34] nuScenes [3]
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Mean cumulative delta error for both ViiNeuS and StreetSurf computed across the four sequences from each dataset as detailed in Tab. 1 of the main paper. Curves indicate the percentage of ground-truth points having an error distance to the closest predicted mesh triangle which is lower than a given value. Light contours represent the standard deviation for each method.
[…]
@@ 10.2 Additional qualitative results
COLMAP [32] OpenMVS [26] GOF [47] StreetSurf [12] ViiNeuS (ours)
@@
COLMAP [32] OpenMVS [26] GOF [47] StreetSurf [12] ViiNeuS (ours)
@@
COLMAP [32] OpenMVS [26] GOF [47] StreetSurf [12] ViiNeuS (ours)
@@
COLMAP [32] OpenMVS [26] GOF [47] StreetSurf [12] ViiNeuS (ours)
@@
Figure 8: Qualitative experiments results on 8(a) KITTI-360s [21], 8(b) Pandaset [43], 8(c) nuScenes [3] and 8(d) Waymo Open Dataset [34]. We compare our mesh extracted from our SDF to GOF, COLMAP, OpenMVS and StreetSurf meshes.
We report in Fig.8 additional qualitative results on nuScenes [3] and Waymo Open Dataset [34]. We find that ViiNeuS reconstructs higher-quality surfaces compared to StreetSurf and can recover many scene details (see highlighted red-boxes on the figures).
@@ 10.3 Ablation study
We ablate the effect of random sample attribution and our proposed probability-based samples attribution. The qualitative results at various training steps are presented in  Fig.9. The results demonstrate that ViiNeuS samples attribution strategy initially learns the coarse geometry of the scene during early training stages. While random sample attribution can approximate an accurate SDF representation by the end of the hybrid stage, it results in an incomplete mesh compared to the mesh generated using probability-based sample attribution.
@@ 11.1 Textured mesh
Due to the high-quality of ViiNeuS’s reconstructed surfaces, we can leverage modern Multi-View Stereo (MVS) tools like OpenMVS [26] to produce detailed and colorized representations of driving sequences. As shown in Fig. 10, we find that ViiNeuS’s textured mesh is more complete and accurate compared to StreetSurf’s textured mesh.
ViiNeuS (ours) Refer to caption Refer to caption
@@
(c) ViiNeuS (ours) (d) StreetSurf [12]
Figure 11: Failure case of ViiNeuS for the sequence 31 from KITTI-360 [21]. We report the ground-truth RGB image and the predicted sky mask. In a different point-of-view than the GT RGB, we compare our generated SDF mesh to StreetSurf’s mesh.
@@
(c) ViiNeuS (ours) (d) StreetSurf [12]
Figure 12: Failure case of ViiNeuS for the sequence 23 from Pandaset [43]. We report the ground-truth RGB image and the wide part of the sequence highlighted. In a different point-of-view than the GT RGB, we compare our generated SDF mesh to StreetSurf’s mesh.
@@
(c) ViiNeuS (ours) (d) StreetSurf [12]
Figure 13: Failure case of ViiNeuS for the sequence 664 from nuScenes [3]. We report the ground-truth RGB image and the monocular normal. In a different point-of-view than the GT RGB we compare our generated SDF mesh to StreetSurf’s mesh.
@@ 12 Limitations
Although ViiNeuS’s reconstructed surfaces are highly detailed and accurate, we find that our method can fail in three distinct scenarios:
  • Disentangling fine details from the sky: unlike StreetSurf [12], which models close range, far range, and sky separately, ViiNeuS separates only the sky from the other scene’s modeling. However, ViiNeuS may struggle to distinguish fine details from the sky, particularly in cases where objects are thin, as shown in Fig. 11.
  • Fine details in wide sequences: ViiNeuS sometimes fails in accurately reconstructing scene details in wide and open sequences, such as scene 23 from Pandaset reported in Fig. 12.
  • Inaccurate road reconstruction: for sequence 664 from nuScenes (see Fig. 13), ViiNeuS faces challenges in reconstructing the road due to inaccuracies in the monocular normal prediction from Omnidata [8]. In comparison, StreetSurf [12] demonstrates more accurate road reconstruction, attributed to its road-surface initialization.
13 Supplementary video
We show in the supplementary video ViiNeuS’s meshes compared to GOF [47] and StreetSurf [12] on one sequence from each of the four evaluated datasets. All meshes were visualized with Blender [6] by animating the camera trajectory to generate the videos. GOF [47] meshes are incomplete at the beginning of scenes and very noisy, as the method is designed for landmark reconstruction and does not address the challenges of driving sequences, such as low image overlap, off-centered regions of interest, the need to handle both close and far-range objects across a wide range of distances, and sky modeling. In addition, in the sky region and empty spaces that are commonly found in driving scenes, GOF tends to create triangles from noisy Gaussians. While the explicit 3DGS formulation is tailored for sparse scenes, it cannot effectively manage the inherent complexities introduced by the specific sensor configurations in driving sequences.